OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1)and perpendicular to y = 1/2x + 3.

OpenStudy (hoblos):

the product of slopes of two perpendicular lines is (-1) use this to get the slope of the line

OpenStudy (anonymous):

soo x is -1?

OpenStudy (hoblos):

m *(1/2) =-1 so m=?

OpenStudy (anonymous):

y=-2x+5 this is what i got

OpenStudy (hoblos):

the slope is -2 .. that is right but when you apply the formula you get y+1 = -2(x+3) y+1 = -2x -6 y= -2x -5 so minus instead of plus

OpenStudy (anonymous):

ohh ok thank you :D

OpenStudy (hoblos):

glad to help :D